# DESY 24-073 May 2024 CERN-TH-2024-068 ZU-TH 26/24

### v1

@@ I Introduction
Drell-Yan lepton pair production processes are among the most important hard scattering events at the LHC. The measured final state contains only leptons. As a result, the corresponding cross sections are known experimentally very precisely. For example, the transverse momentum distribution of Drell-Yan lepton pairs reaches a precision of 0.2 % for the normalized spectra at low values of $p_{T}^{(\ell\ell)}$ ATLAS:2019zci ; CMS:2018mdl . The importance of the process is shown by its frequent use in precision extraction of the parameters in the Standard Model, such as parton distribution functions (PDFs) and the strong coupling constant Alekhin:2017kpj ; Bailey:2020ooq ; Ball:2018iqk . It is also used in the determination of the mass of the $W^{\pm}$-boson ATLAS:2017rzl ; LHCb:2021bjt . All these measurements require a reduction of the theoretical uncertainties to match the experimental ones. The current state of the art in the theory description has advanced significantly in recent years. It has reached next-to-next-to-next-to-leading order (N3LO) accuracy at fixed order in quantum chromodynamics (QCD) perturbation theory Duhr:2020sdp ; Duhr:2021vwj at the inclusive level. Transverse momentum resummation is known at the next-to-next-to-next-to-leading logarithmic (N3LL) level Camarda:2021ict ; Re:2021con ; Chen:2022cgv and even at approximate N4LL Neumann:2022lft ; Camarda:2023dqn accuracy for differential distributions in perturbation theory. At present, the fully differential calculations at N3LO accuracy employ transverse momentum subtractions Catani:2007vq that neglect power corrections and necessarily rely on predictions for vector boson + jet production at the N2LO accuracy. Hence, the reliable computation of cross sections at N2LO in QCD is a prerequisite for further developments in the field.
@@
The publication triggered discussions among the authors of the relevant codes, which resulted in a better understanding of the neglected terms and improvements in the computations. In this paper we provide an update of the comparisons carried out in Ref. Alekhin:2021xcu . The authors of the codes have provided new predictions for the benchmark calculations that we present in Sec. II, now showing excellent agreement. Based on this validation of fixed-order perturbative QCD calculations through NNLO we then study the impact of fiducial cuts on the decay leptons. To that end, we focus on the case of symmetric cuts on the transverse momenta of the leptons, as they are routinely imposed in experimental analyses but display certain unphysical features Frixione:1997ks . In Sec. III, we study cuts on the transverse momenta staggered in a range from a few tens of MeV to a few GeV in perturbative QCD, both at fixed order and applying all-orders resummation of large lepton-pair transverse momentum logarithms. We comment on proposed modifications of the fiducial cuts put forward recently Salam:2021tbm . Since the pathology observed in fixed-order perturbation theory arises from the region of small lepton-pair transverse momentum we can cure it with resummation of the small $p_{T}$ region. We do so in Section IV and find only small differences between the resummed results and the available fixed-order codes. We provide a short discussion of the experimental resolution to gauge the impact of our findings on current analyses in Section V. We summarize in Sec. VI and conclude with our comments on the choice of cuts for experimental Drell-Yan analyses. Information on the set-up and input parameters for some of the codes used for the benchmark comparisons is listed in the Appendices A-D.
@@ II.1 Set-up and code validation
The initial benchmark Alekhin:2021xcu also considered MCFM (version 9.0) Campbell:2019dru and found large effects from linear power corrections in comparison to FEWZ results (local subtraction). The MCFM code [^6] Code available from https://mcfm.fnal.gov/. implements the NNLO computation of Ref. Boughezal:2016wmq and applies $N$-jettiness subtraction Boughezal:2015dva ; Gaunt:2015pea with $\tau_{\rm cut}$ as the jettiness slicing parameter. Since the recoil prescription Catani:2015vma ; Grazzini:2015wpa ; Ebert:2020dfc to remove the linear power corrections cannot be easily adapted to $N$-jettiness slicing and no relevant update of MCFM is available, we refer to Ref. Alekhin:2021xcu for predictions with this code.
@@ II.2 NNLO benchmark predictions
Figure 1: The pulls for the ATLAS data measured in inclusive $pp\to W^{\pm}+X\to l^{\pm}\nu+X$ and $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV Aaboud:2016btc with the statistical (inner bar) and the total uncertainties, including the systematic ones. The fiducial cuts on the decay leptons in the final state are indicated in the figure. The ABMP16 central predictions at NNLO are obtained with FEWZ and the deviations of the predictions from NNLOJET are shown (dashed) for comparison.
@@
The findings of this benchmark study are summarized Fig. 4, which compares in each case the best predictions of DYTURBO, MATRIX (v2.1) and NNLOJET. The predictions of MATRIX (v2.1) and NNLOJET are in excellent agreement, with deviations being often less than one per-mille, i.e. the target uncertainty from the numerical integration. Also the DYTURBO predictions agree very well, typically within two per-mille with the NNLOJET ones, except for the $Z$-boson production with forward leptons, where agreement is at the level of a few per-mille. Fig. 4 also shows, that the predictions from DYTURBO, MATRIX (v2.1) and NNLOJET are all aligned, especially for the case of $W^{\pm}$-boson production. The normalization to the FEWZ predictions introduces some fluctuations, which are due to the numerical integration uncertainties in the FEWZ results, being at the level of a few per-mille.
The DØ data on the electron charge asymmetry distribution $A_{e}$ has been obtained as a function of the electron pseudo-rapidity from $W^{\pm}$-boson production at $\sqrt{s}=1.96$ TeV at the Tevatron D0:2014kma . Theoretical predictions for the benchmark studies are numerically challenging due to large cancellations in the asymmetry. With predictions from NNLOJET and the update of MATRIX (version 2.1) we are in a position to check the FEWZ results, already presented in Ref. Alekhin:2021xcu . This is illustrated in Fig. 5, where we plot the NNLO predictions obtained with those codes. The numbers obtained with NNLOJET and MATRIX (version 2.1) are in excellent agreement and are also compatible with the FEWZ predictions within the substantially larger uncertainties from the numerical integration of the latter. The numerical integration uncertainties of the NNLOJET and MATRIX numbers are negligible on the scale of Fig. 5, while those of FEWZ are indicated in the plot.
@@ III Theoretical formalism for lepton fiducial cuts
III.1 Definition of staggered fiducial cuts
@@
Figure 6: The cross sections for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV at LO (dotted), NLO (dashed) and NNLO (solid) in QCD with ABMP16 PDFs and $y_{ll}\leq 2.5$ computed with NNLOJET as a function of $\Delta p_{T}\in[-5,5]$ GeV defined in Eq. (3) for the staggered fiducial cuts on the decay leptons in the final state (left plot) and zoom on the range $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot).
@@
We note that in the literature staggered cuts have also been defined by applying cuts on identified leptons (electron and positron in neutral-current Drell-Yan or lepton and neutrino in charged current Drell-Yan). In this case it can be shown Grazzini:2017mhc ; Alekhin:2021xcu ; Buonocore:2021tke that the performance of slicing methods are much improved compared to symmetric cuts or to staggered cuts imposed on leading and subleading leptons, as they lead to a quadratic sensitivity in the region $q_{T}<\Delta p_{T}$. However, we expect a similar behavior of the fixed-order cross section as a function of $\Delta p_{T}$ as that discussed for our definition of staggered cuts.
@@
To investigate the impact of these considerations on the cross section for $Z$ production at $\sqrt{s}=7$ TeV with a cut on the lepton pseudorapidities $|\eta_{\ell_{1},\ell_{2}}|\leq 2.5$ as a function of $\Delta p_{T}$, we present the fixed-order predictions of NNLOJET in Fig. 6 and MATRIX in Fig. 7. Results at LO, NLO and NNLO in perturbative QCD for the staggered lepton decay cuts defined in Eq. (3) are shown. Excellent agreement between the NNLOJET and the MATRIX (v2.1) numbers is observed for all values of $\Delta p_{T}$. The latter have been obtained accounting for the linear power corrections in the $q_{T}$-subtraction and bin-wise $r_{\rm cut}$ extrapolation, as discussed previously. The logarithmic divergences described above lead to the kink at $\Delta p_{T}=0$ and the resulting non-monotonic behavior as $\Delta p_{T}$ is varied in these plots. This result is a pathology of fixed-order perturbation theory when symmetric cuts are applied.
@@ III.3 Definition of product cuts
Figure 8: Same as in Fig. (6), now with product cuts on the transverse momenta of the decay leptons in the final state. The product cuts are defined in Eq. (12) and ranges $\Delta p_{T}\in[-5,5]$ GeV (left plot) and $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot) are shown.
Figure 9: The ratio of cross sections for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production from Fig. 8 with the staggered product cuts of Eq. (12) over those from Fig. 6 with linear asymmetric cuts of Eq. (3) as a function of $\Delta p_{T}$ at NLO (dashed) and NNLO (solid) in QCD. The ranges $\Delta p_{T}\in[-5,5]$ GeV (left plot) and $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot) are shown.
@@
The perturbative expansion for the cross section in this case depends only quadratically on the $Z$-boson transverse momentum, and the unphysical behavior as $\Delta p_{T}\to 0$ is therefore expected to decrease Salam:2021tbm . We show the predictions for these product cuts using NNLOJET in Fig. 8. A comparison of the relative difference between the regular cuts and the product cuts at NLO and NNLO is shown in Fig. 9. While the unphysical dependence on $\Delta p_{T}$ is lessened by switching to product cuts it is still clearly visible on the plots. We note that both the fixed-order and all-orders structure of the cross section with product cuts defined in Eq. (12) are analogous to those discussed in the previous subsections, including the possibility of a mismatch between linear slopes due to the way $\Delta p_{T}$ is defined here.

### v2

@@ I Introduction
Drell-Yan lepton pair production processes are among the most important hard scattering events at the LHC. The measured final state contains only leptons. As a result, the corresponding cross sections are known experimentally very precisely. For example, the transverse momentum distribution of Drell-Yan lepton pairs reaches a precision of 0.2 % for the normalized spectra at low values of $p_{T}^{(\ell\ell)}$ ATLAS:2019zci ; CMS:2018mdl . The importance of the process is shown by its frequent use in precision extraction of the parameters in the Standard Model, such as parton distribution functions (PDFs) and the strong coupling constant Alekhin:2017kpj ; Bailey:2020ooq ; Ball:2018iqk . It is also used in the determination of the mass of the $W^{\pm}$-boson ATLAS:2017rzl ; LHCb:2021bjt . All these measurements require a reduction of the theoretical uncertainties to match the experimental ones. The current state of the art in the theory description has advanced significantly in recent years. It has reached next-to-next-to-next-to-leading order (N3LO) accuracy at fixed order in quantum chromodynamics (QCD) perturbation theory Duhr:2020sdp ; Duhr:2021vwj at the inclusive level. Transverse momentum resummation is known at the next-to-next-to-next-to-leading logarithmic (N3LL) level Camarda:2021ict ; Re:2021con ; Ju:2021lah ; Chen:2022cgv and even at approximate N4LL Neumann:2022lft ; Camarda:2023dqn accuracy for differential distributions in perturbation theory. At present, the fully differential calculations at N3LO accuracy employ transverse momentum subtractions Catani:2007vq that neglect power corrections and necessarily rely on predictions for vector boson + jet production at the N2LO accuracy. Hence, the reliable computation of cross sections at N2LO in QCD is a prerequisite for further developments in the field.
@@
The publication triggered discussions among the authors of the relevant codes, which resulted in a better understanding of the neglected terms and improvements in the computations. In this paper we provide an update of the comparisons carried out in Ref. Alekhin:2021xcu . The authors of the codes have provided new predictions for the benchmark calculations that we present in Sec. II, now showing excellent agreement. Based on this validation of fixed-order perturbative QCD calculations through NNLO we then study the impact of fiducial cuts on the decay leptons. To that end, we focus on the case of symmetric cuts on the transverse momenta of the leptons, as they are routinely imposed in experimental analyses but display certain unphysical features Frixione:1997ks . In Sec. III, we study cuts on the transverse momenta staggered in a range from a few tens of MeV to a few GeV in perturbative QCD, both at fixed order and applying all-orders resummation of large lepton-pair transverse momentum logarithms. We comment on proposed modifications of the fiducial cuts put forward recently Salam:2021tbm . Since the pathology observed in fixed-order perturbation theory arises from the region of small lepton-pair transverse momentum we can cure it with resummation of the small $p_{T}$ region. We do so in Section IV and find only small differences between the resummed results and the available fixed-order codes. We provide a short discussion of the experimental resolution to gauge the impact of our findings on current analyses in Section V. We summarize in Sec. VI and conclude with our comments on the choice of cuts for experimental Drell-Yan analyses. Cross sections and information on the set-up and input parameters for some of the codes used for the benchmark comparisons are listed in the Appendices AE.
@@ II.1 Set-up and code validation
The initial benchmark Alekhin:2021xcu also considered MCFM (version 9.0) Campbell:2019dru and found large effects from linear power corrections in comparison to FEWZ results (local subtraction). The MCFM code [^6] Code available from https://mcfm.fnal.gov/. implements the NNLO computation of Ref. Boughezal:2016wmq and applies $N$-jettiness subtraction Boughezal:2015dva ; Gaunt:2015pea with $\tau_{\rm cut}$ as the jettiness slicing parameter. Since the recoil prescription Catani:2015vma ; Grazzini:2015wpa ; Ebert:2020dfc to remove the linear power corrections cannot be easily adapted to $N$-jettiness slicing we refer to Ref. Alekhin:2021xcu for predictions with this version of the code. As of MCFM (version 10.0) Neumann:2022lft the code also allows for all NNLO calculations implemented (as well as for the N3LO calculations for charged- and neutral-current Drell-Yan) to be performed using $q_{T}$-subtraction, also accounting for fiducial power corrections, see, e.g. Ref. Campbell:2023lcy ; Campbell:2024hjq .
@@ II.2 NNLO benchmark predictions
Figure 1: Relative deviation of ATLAS data measured in inclusive $pp\to W^{\pm}+X\to l^{\pm}\nu+X$ and $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV Aaboud:2016btc with the statistical (inner bar) and the total uncertainties, including the systematic ones. The fiducial cuts on the decay leptons in the final state are indicated in the figure. The ABMP16 central predictions at NNLO are obtained with FEWZ and the deviations of the predictions from NNLOJET are shown (dashed) for comparison.
@@
The findings of this benchmark study are summarized Fig. 4, which compares in each case the best predictions of DYTURBO, MATRIX (v2.1) and NNLOJET. The predictions of MATRIX (v2.1) and NNLOJET are in excellent agreement, with deviations being often less than one per-mille, i.e. the target uncertainty from the numerical integration. Also the DYTURBO predictions agree very well, typically within two per-mille with the NNLOJET ones, except for the $Z$-boson production with forward leptons, where agreement is at the level of a few per-mille. Fig. 4 also shows, that the predictions from DYTURBO, MATRIX (v2.1) and NNLOJET are all aligned, especially for the case of $W^{\pm}$-boson production. The normalization to the FEWZ predictions introduces some fluctuations, which are due to the numerical integration uncertainties in the FEWZ results, being at the level of a few per-mille. The predictions obtained with NNLOJET are listed in App. A.
The DØ data on the electron charge asymmetry distribution $A_{e}$ has been obtained as a function of the electron pseudo-rapidity from $W^{\pm}$-boson production at $\sqrt{s}=1.96$ TeV at the Tevatron D0:2014kma . Theoretical predictions for the benchmark studies are numerically challenging due to large cancellations in the asymmetry. With predictions from NNLOJET and the update of MATRIX (version 2.1) we are in a position to check the FEWZ results, already presented in Ref. Alekhin:2021xcu . This is illustrated in Fig. 5, where we plot the NNLO predictions obtained with those codes. The numbers obtained with NNLOJET and MATRIX (version 2.1) are in excellent agreement and are also compatible with the FEWZ predictions within the substantially larger uncertainties from the numerical integration of the latter. The numerical integration uncertainties of the NNLOJET and MATRIX numbers are negligible on the scale of Fig. 5, while those of FEWZ are indicated in the plot. All numbers computed with NNLOJET are also given in App. A.
@@ III Theoretical formalism for lepton fiducial cuts
III.1 Definition of linear asymmetric fiducial cuts
@@
Figure 6: The cross sections for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV at LO (dotted), NLO (dashed) and NNLO (solid) in QCD with ABMP16 PDFs and $y_{ll}\leq 2.5$ computed with NNLOJET as a function of $\Delta p_{T}\in[-5,5]$ GeV defined in Eq. (3) for the linear asymmetric fiducial cuts on the decay leptons in the final state (left plot) and zoom on the range $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot).
@@
We note that in the literature so-called staggered cuts have also been defined by applying cuts on identified leptons (electron and positron in neutral-current Drell-Yan or lepton and neutrino in charged current Drell-Yan). In this case it can be shown Grazzini:2017mhc ; Alekhin:2021xcu ; Buonocore:2021tke that the performance of slicing methods are much improved compared to symmetric cuts or to staggered cuts imposed on leading and subleading leptons, as they lead to a quadratic sensitivity in the region $q_{T}<\Delta p_{T}$. However, we expect a similar behavior of the fixed-order cross section as a function of $\Delta p_{T}$ as that discussed for our definition of linear asymmetric cuts.
@@
To investigate the impact of these considerations on the cross section for $Z$ production at $\sqrt{s}=7$ TeV with a cut on the lepton pseudorapidities $|\eta_{\ell_{1},\ell_{2}}|\leq 2.5$ as a function of $\Delta p_{T}$, we present the fixed-order predictions of NNLOJET in Fig. 6 and MATRIX in Fig. 7. Results at LO, NLO and NNLO in perturbative QCD for the linear asymmetric lepton decay cuts defined in Eq. (3) are shown. Excellent agreement between the NNLOJET and the MATRIX (v2.1) numbers is observed for all values of $\Delta p_{T}$. The latter have been obtained accounting for the linear power corrections in the $q_{T}$-subtraction and bin-wise $r_{\rm cut}$ extrapolation, as discussed previously. The logarithmic divergences described above lead to the kink at $\Delta p_{T}=0$ and the resulting non-monotonic behavior as $\Delta p_{T}$ is varied in these plots. This result is a pathology of fixed-order perturbation theory when symmetric cuts are applied.
@@ III.3 Definition of product cuts
Figure 8: Same as in Fig. 6, now with product cuts on the transverse momenta of the decay leptons in the final state. The product cuts are defined in Eq. (12) and ranges $\Delta p_{T}\in[-5,5]$ GeV (left plot) and $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot) are shown.
Figure 9: The ratio of cross sections for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production from Fig. 8 with the product cuts of Eq. (12) over those from Fig. 6 with linear asymmetric cuts of Eq. (3) as a function of $\Delta p_{T}$ at NLO (dashed) and NNLO (solid) in QCD. The ranges $\Delta p_{T}\in[-5,5]$ GeV (left plot) and $\Delta p_{T}\in[-0.3,0.3]$ GeV (right plot) are shown.
@@
The perturbative expansion for the cross section in this case depends only quadratically on the $Z$-boson transverse momentum, and it is interesting to study whether the unphysical behavior as $\Delta p_{T}\to 0$ improves in this case. We show the predictions for these product cuts using NNLOJET in Fig. 8. A comparison of the relative difference between the regular cuts and the product cuts at NLO and NNLO is shown in Fig. 9. While the unphysical dependence on $\Delta p_{T}$ is lessened by switching to product cuts it is still clearly visible on the plots. We note that both the fixed-order and all-orders structure of the cross section with product cuts defined in Eq. (12) are analogous to those discussed in the previous subsections, including the possibility of a mismatch between linear slopes due to the way $\Delta p_{T}$ is defined here.
III.4 Rapidity distributions
Figure 10: Deviation of the rapidity distribution for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV as a function of $\Delta p_{T}$ relative to the results for $\Delta p_{T}=0$ at NNLO in QCD computed with NNLOJET. Shown are selected $\Delta p_{T}$ values (indicated by color) using the definition in Eq. (3) for linear asymmetric cuts (left plot) and the one of Eq. (12) for product cuts (right plot). Solid lines denote negative values of $\Delta p_{T}$ as indicated in the plots, dotted lines of the same color display the corresponding $\Delta p_{T}$ values with positive sign.
Figure 11: The cross sections for $pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV at LO (dotted), NLO (dashed) and NNLO (solid) in QCD with ABMP16 PDFs computed with NNLOJET as a function of $\Delta p_{T}\in[-5,5]$ GeV defined in Eq. (3) for the linear asymmetric fiducial cuts on the decay leptons in the final state. The rapidity bins $|y_{ll}|\in[0.4,0.6]$ (left plot) and $|y_{ll}|\in[1.4,1.6]$ (right plot) are shown.
Figure 12: Fiducial NNLO to NLO $K$-factor as a function of $|y_{\ell\ell}|$ for linear asymmetric cuts (left) and product cuts (right) for different values of $\Delta p_{T}$, normalised to the inclusive NNLO to NLO $K$-factor.
It is instructive to revisit the study of the $\Delta p_{T}$ dependence of the cross section for $Z$-boson production with central leptons, which directly relates to the numerical studies presented in Sec. II. For selected fixed values of $\Delta p_{T}$ we display in Fig. 10 the $K$-factor $\sigma_{\rm NNLO}/\sigma_{\rm NLO}$ relative to the inclusive case for asymmetric cuts (left) and product cuts (right) as a function of the $Z$-boson rapidity. In both figures we also display the $K$-factor with symmetric cuts in black.
For the selected values of $\Delta p_{T}$, we observe a difference at the 2% level across the whole range in rapidity between the asymmetric case and the inclusive case. There is a turning point halfway through the rapidity range considered, with the ratio moving from $\simeq-2\%$ at central rapidity to $\simeq+2\%$ at $|y_{\ell\ell}|=2.2$, with an abrupt change in slope starting around $|y_{\ell\ell}|\simeq 1$. The symmetric cuts case displays a similar trend at large $|y_{\ell\ell}|$, while it is closer to the inclusive case at central rapidities. Also in this case, we observe sudden changes in slope at central rapidities, even though the overall deviation with respect to the inclusive case is between [-0.5%, +0.5%] for $|y_{\ell\ell}|$ below 1.5.
On the other hand, the product cut curves display a smoother behaviour, and they are almost indistinguishable from the inclusive result for $|y_{\ell\ell}|$ below 1.2. Above this value, the $K$-factors start to depart from the inclusive case with a slope similar to the one observed in the asymmetric and the symmetric case, with differences up to $\simeq+2\%$. For the kinematics under consideration ($pp\to Z/\gamma^{*}+X\to l^{+}l^{-}+X$ production at $\sqrt{s}=7$ TeV with cuts on the decay leptons $p_{T}^{\ell}\geq 20\,\text{GeV}$), linear power corrections in $\Delta p_{T}$ appear for central rapidities, $y_{ll}\lesssim 1.0$, due to the fiducial cuts breaking the azimuthal symmetry in the integral over the lepton decay phase space, as discussed in Refs. Ebert:2019zkb , Ebert:2020dfc . For larger rapidities $y_{ll}\gtrsim 1.5$ the rapidity constraints dominate the integral over the lepton phase space and azimuthal symmetry is restored, resulting in quadratic power corrections in $\Delta p_{T}$. This leads to only small deviations of all predictions from the case of symmetric cuts. The impact of the cuts on the lepton decay phase space has been illustrated in our previous study Alekhin:2021xcu , including the region around around $y_{ll}\simeq 1.2$, where the transition occurs 777 The impact of the fiducial cuts applied to $Z/\gamma^{*}$-boson production for different values of the gauge-boson rapidity $y_{ll}$ on the real emission phase space of the decay leptons has been illustrated in Fig. 15 of Ref. Alekhin:2021xcu ..
The fact that the product cut $K$-factors display a better behaviour at central rapidities shows that, also at the differential level, product cuts are free of linear power corrections, which are instead responsible for the larger differences observed in the asymmetric case and, to some extent, in the symmetric case. The effect observed are, however, limited to a few percent, and do not display the much larger differences observed in the Higgs case when asymmetric cuts are enforced Chen:2021isd . Moreover, our analysis is performed at NNLO, while larger shape distortions were observed in the Higgs case at N3LO. A better assessment of the impact of linear power corrections associated with the choice of cuts requires to consider resummation effects at NNLL and beyond, which we will discuss in the next section.